Complete the following table. ​(Round your answers to the nearest whole​ number.)

More Info Compound [F] Eutectic Temperature [C] [K] 655 [OR] (a) 5% Sn, 95% Zn 48% Sn, 52% In (b) 704
Answer #1

First let me write some relations

F = 1.8C + 32 (i)

K = C + 273.15 (ii)

R= F + 459.67 (iii)

where F , C , R and K are temperature in fahrenheit , celcius , rankine a) 655k K=C + 273,15 °C 2 655-273.15 T +381-85°C of 1.8 (°C) +32 2 1.8 (381.85) +32 - 719.33°F •R = of + 459.67 = 719.33 4 4